The consumer utility function for mango tree and orange tree is given U=q1q2, the price for mango tree is $40.00 and orange is $20.00. The consumer income for the period is $120.00. Determine the quantities of mango tree and orange tree which should be purchased in order to maximize derived utility?.


1
Expert's answer
2021-05-03T10:54:50-0400

Solution:

The optimal consumption bundle is one where the slope of the indifference curve "\frac{MUq1}{MUq2}" is equal to the slope of the budget line.


Derive MRSq1q2 = "\frac{MUq1}{MUq2}"

MUq1 = ​"\frac{∂U}{∂q1}" = q2


MUq2 = ​"\frac{∂U}{∂q2}" = q1


MRSq1q2 = "\frac{q_{2} }{q_{1} }"


MRSq1q2 = "\frac{Pq1}{Pq2}"


"\frac{q_{2} }{q_{1} }" = "\frac{40 }{20 }"

q2 = 2q1

Derive utility function:

M = Pq1q1+ Pq2q2

120 = 40q1 + 20q2

Therefore:

120 = 40q1 + 20(2q1)

120 = 40q1 + 40q1

120 = 80q1

q1 = "\frac{120}{80 }" = 1.5


q1 = 1.5

Plug this figure into the q2 function:

q2 = 2q1

q2 = 2(1.5)

q2 = 3

Maximizing Utility (1.5, 3)

The quantities of the mango tree to be purchased to maximize derived utility = 1.5

The quantities of the orange tree to be purchased to maximize derived utility = 3